OpenStudy (anonymous):

Anti-derivative of (1/t^(2)-1/t^(4)) ?

jimthompson5910 (jim_thompson5910):

I'll get you started \[\Large \int\left(\frac{1}{t^2}-\frac{1}{t^4}\right)dt\] \[\Large \int\left(\frac{1}{t^2}\right)dt-\int\left(\frac{1}{t^4}\right)dt\] \[\Large \int\left(t^{-2}\right)dt-\int\left(t^{-4}\right)dt\]

OpenStudy (anonymous):

ok I got -t^(-1)-((-1/3)t^(-3)) is that right?

jimthompson5910 (jim_thompson5910):

correct, \[\Large \int\left(\frac{1}{t^2}-\frac{1}{t^4}\right)dt\] \[\Large \int\left(\frac{1}{t^2}\right)dt-\int\left(\frac{1}{t^4}\right)dt\] \[\Large \int\left(t^{-2}\right)dt-\int\left(t^{-4}\right)dt\] \[\Large -t^{-1} +\frac{1}{3}t^{-3}+C\] \[\Large -\frac{1}{t}+ \frac{1}{3t^{3}}+C\]

OpenStudy (anonymous):

OK I got it now. Thank you so much :)

jimthompson5910 (jim_thompson5910):

you're welcome